# Universal 3-Edge-Weightability of Regular Graphs

### v1

@@ 2. Proof of Theorem 1
After resolving conflicts for $X_{2^{+}}$, we update $I_{0,0}$. Note that $X_{1}$ remains unchanged. Let $X^{\ast}_{1}\subseteq X_{1}$ be the set of vertices in the current $X_{1}$ that have a neighbor in (updated) $I_{0,0}$. If $X^{\ast}_{1}=\emptyset$, we are finished.
Assume $X^{\ast}_{1}\neq\emptyset$ and let $M^{\ast}_{1}$ be the subset of $M_{1}$ consisting of edges incident to vertices in $X^{\ast}_{1}$. Recall that each edge in $M^{\ast}_{1}$ connects a vertex $x\in X^{\ast}_{1}$ to a Type II vertex $z\in\cup_{i\geq 2}I_{i}$, and each edge in $M^{\ast}_{1}$ is unweighted. Assign weight $d_{1}$ to every edge in $M^{\ast}_{1}$. After the assignment:
  • $d_{w}(x)$ changes from 0 to $d_{1}$;
  • $d_{w}(y_{x})$ remains 0 (since edge $xy_{x}$ is not in $M^{\ast}_{1}$ and keeps weight 0);
  • $d_{w}(z)$ decreases by $d_{1}$ for the corresponding $z\in\cup_{i\geq 2}I_{i}$ which becomes Type III.
This implies that condition (5) holds, and (1), (2), (3) remain valid. Note that, at present, both $X^{\ast}_{2^{+}}$ and $X^{\ast}_{1}$ are empty, and so (6) holds. Condition (4) holds because: vertices with $d_{w}(x)=-3d_{1}$ come from Case 2 or Case 3. In Case 2, $x\in X_{3^{+}}$ has at least three neighbors in $I_{0}$; in Case 3, $x\in X_{2}$ has exactly two neighbors in $I_{0}$. In both cases, all neighbors $y$ of $x$ in $I_{0}$ satisfy $d_{w}(y)\neq-3d_{1}$ by construction. Thus, $w$ is a proper $\{-d_{1},0,d_{2}\}$-edge weighting of $G$. This completes the proof of Theorem 1. ∎
@@ 3. Concluding Remarks
Theorem 1 and Corollary 2 establish that every regular graph is universally 3-edge-weightable. The independent-set decomposition technique, originally developed for total weightings, proves adaptable to edge-weightings with arbitrary 3-element set of weights.
Our work suggests that regularity provides a structural framework within which arbitrary three-element weight set can be accommodated. The conflict-resolution scheme developed here may find applications in other weighting and labeling problems. The two-phase approach—global matching adjustments followed by local conflict resolution—offers a template for handling parameter-dependent constructions in combinatorial graph theory.
Future work. Several natural questions arise from our result:

### v2

@@ 2. Proof of Theorem 1
After resolving conflicts for $X_{2^{+}}$, we update $I_{0,0}$. Note that $X_{1}$ remains unchanged. Let $X^{\ast}_{1}\subseteq X_{1}$ be the set of vertices in $X_{1}$ that have a neighbor in (updated) $I_{0,0}$. If $X^{\ast}_{1}=\emptyset$, we are finished.
Assume $X^{\ast}_{1}\neq\emptyset$ and let $M^{\ast}_{1}$ be the subset of $M_{1}$ consisting of edges incident to vertices in $X^{\ast}_{1}$. Assign weight $d_{1}$ to every edge in $M^{\ast}_{1}$. After the assignment, for each $x\in X^{\ast}_{1}$, $d_{w}(x)=-d_{1}$, while $d_{w}(y_{x})$ remains 0 and $d_{w}(z_{x})$ decreases by $d_{1}$, making $z_{x}$ a Type III vertex. This implies that condition (5) holds, and conditions (1)-(3) remain valid. Note that, at present, both $X^{\ast}_{2^{+}}$ and $X^{\ast}_{1}$ are empty, and so condition (6) holds. Condition (4) holds because: vertices with $d_{w}(x)=-3d_{1}$ come from Case 2. In Case 2.1, $x\in X_{3^{+}}$ has at least three neighbors in $I_{0}$; in Case 2.2, $x\in X_{2}$ has exactly two neighbors in $I_{0}$. In both cases, all neighbors $y$ of $x$ in $I_{0}$ satisfy $d_{w}(y)\neq-3d_{1}$ by construction. Thus, conditions (1)-(6) all hold, and $w$ is a proper $\{-d_{1},0,d_{2}\}$-edge weighting of $G$.
This completes the proof of Theorem 1. ∎
@@ 3. Concluding Remarks
This paper establishes that every regular graph is universally 3-edge-weightable. The conflict-resolution scheme introduced in this work may be applicable to other weighting and labeling problems. The two-phase approach–comprising global matching adjustments followed by local conflict resolution–provides a framework for handling parameter-dependent constructions in combinatorial graph theory.
Future work. Our result naturally leads to several open questions: